# Capturing Knowledge Graphs and Rules with Octagon Embeddings

### v1

@@ 1 Introduction
To address these concerns, our aim in this paper is to develop a region based model which is as simple as possible, while (i) still being expressive enough to capture arbitrary knowledge graphs and (ii) using regions which are closed under intersection and composition. Based on these desiderata, we arrive at a model which relies on axis-aligned octagons (with all angles fixed at 45 degrees). We show that despite their simplicity, the proposed octagon embeddings are sufficiently expressive to properly capture a large class of rule bases. This is an important property, among others because it means that we can inject prior knowledge, in the form of a given rule base, without unintended consequences. Our result is considerably more general than what is possible with BoxE, which is not able to capture any closed path rules, and more general than what is known about ExpressivE. Moreover, because compositions and intersections of octagons can be straightforwardly computed, octagon embeddings are considerably more practical than existing alternatives. While our main focus is on better understanding the expressivity of region based models, we have also empirically evaluated the proposed octagon embeddings. We found octagons to achieve results close to the current state-of-the-art. This demonstrates that learning octagon embeddings is a promising strategy, especially in contexts where both knowledge graphs and rules need to be modelled.
@@ 2 Preliminaries
(a) TransE bands
(b) Hexagons
Figure 1: Basic region based embeddings.
@@ 4.2 Capturing Rules
[…]
@@
[…]
@@ 5 Learning Octagon Embeddings
where we write $\mathbf{u}_{c}=(\mathbf{u}^{+}+\mathbf{u}^{-})/2$ and $\mathbf{u}_{w}=(\mathbf{u}^{+}-\mathbf{u}^{-})/2$. The functions $\text{dist}_{x}$, $\text{dist}_{y}$ and $\text{dist}_{v}$ are defined analogously. Note that $\mathbf{e}\oplus\mathbf{f}\in X_{r}$ iff $\text{dist}_{x}(e,r,f)\geq 0.5$, $\text{score}_{y}(e,r,f)\geq 0.5$, $\text{dist}_{u}(e,r,f)\geq 0.5$ and $\text{dist}_{v}(e,r,f)\geq 0.5$. Furthermore note that BoxE and ExpressivE also convert interval constraints into soft scores. However, rather than using scores of the form (17), these models use a piecewise linear function of the distance to the centre of the interval. Although a different slope is used for points inside and outside the interval, we may question to what extent the resulting embeddings still define regions. While the regions defined by scoring functions of the form (17) also have fuzzy boundaries, this “fuzziness” is mostly limited to the immediate vicinity of the boundary.
@@
where we write $\odot$ for component-wise multiplication. Weighted variants of $\text{dist}_{x}$, $\text{dist}_{y}$ and $\text{dist}_{v}$ are defined analogously. The attention weights intuitively allow the model to “forget” certain constraints, e.g. by focusing only (or primarily) on the $u$ band in some coordinates. An overall score is calculated as follows (and similar for the weighted variant):
$s(e,r,f)=-\|\sum_{i\in\{u,v,x,y\}}\text{dist}_{i}(e,r,f)\|_{p}$ (19)
As in BoxE and ExpressivE, coordinate-wise scores are aggregated to a scalar by taking their negated norm. In the above expression, $p$ is a hyperparameter. Learning occurs by minimizing the following margin loss function with self-adversarial negative sampling Sun et al. (2019):
@@
where $\mathbf{\alpha}=\text{softmax}(\left[s(e_{1},r,f_{1});\ldots;s(e_{n},r,f_{n})\right]$, the margin $\lambda$ is a hyperparameter, and $N=\{(e_{1},r,f_{1}),\allowbreak\ldots,\allowbreak(e_{n},r,f_{n})\}$ is a set of randomly sampled negative examples.
@@ 6 Experimental Results
We consider several variants of our model. The full model is denoted by uvxy, whereas ux refers to a model in which $\text{dist}_{v}$ and $\text{dist}_{y}$ are not used, and similar for the other variants. For $uv$ and $uvxy$, we have used the variant with attention weights (18). For the other configurations, attention weights were not used, as they were found to make little difference [^3] A more complete analysis of the impact of attention weights can be found in the appendix.. Note that our $u$ model is almost identical to TransE, except that an explicit width parameter is used, while uv can be seen as a variant of ExpressivE without learnable scale factors.
@@
$u$ 22.8 37.2 53.2 33.0 00.4 40.4 52.0 21.6
$ux$ 22.9 36.4 52.6 32.7 00.4 40.4 53.2 22.0
$uxy$ 22.8 36.5 52.3 32.6 01.7 46.7 53.8 22.5
$uv^{*}$ 23.0 35.7 51.1 32.3 42.8 46.5 50.1 45.5
$uvxy^{*}$ 22.8 35.6 50.6 32.0 44.3 49.9 56.6 48.5
Table 1: Link prediction performances of region based embedding models. Configurations with $*{}^{*}$ use the variant with attention weights.
@@
The performance of $u$ is almost identical to TransE, which suggests that the addition of an explicit width parameter plays no significant role. Adding the $x$, $y$ and $v$ constraints does not bring any benefits on FB15k-237. On WN18RR, however, introducing the $v$ constraint has a big impact on H@1, which goes from less than 1% for $u$, $ux$ and $uxy$ to more than 40% for $uv$ and $uvxy$. This result is in line with Proposition 5, which states that hexagons cannot capture relations that are both symmetric and irreflexive: for most WN18RR relations, models without $v$ constraints incorrectly rank $(e,r,e)$ triples first. Including the $x$ and $y$ constraints also clearly helps the model in the case of WN18RR. Introducing attention weights, as in (18), has a significant positive effect on the overall performance on WN18RR but not for FB15k-237. For instance, without attention weights, $uvxy$ achieves a H@1 of 4.8, H@3 of 43.5, H@10 of 53.4 and MRR of 25.5 (not shown). As Table 1 shows, $uvxy$ outperforms BoxE on all metrics. Further analysis can be found in the appendix.
Overall, we find that octagon embeddings slightly, but consistently, underperform ExpressivE. We may wonder whether this reflects the inherent limitations of using fixed angles. In fact, when inspecting the scale factors in learned ExpressivE embedddings, we noticed that most are close to $-1$, 00 or $1$ (see Figure 4). In other words, most of the learned parallelograms are diamonds ($uv$), bands ($u$, $v$, $x$, $y$) and squares ($xy$), all of which can be represented in our octagon model. This suggests that the underperformance of the octagon model is rather the result of overparameterisation. On the one hand, we have shown theoretically that all four constraints are needed to capture semantic dependencies, and this is also confirmed by our empirical results on WN18RR. On the other hand, not all constraints might be needed for all coordinates. The use of attention weights was inspired by this view, but the weights themselves of course also introduce further parameters.
@@ 7 Related Work
Knowledge graph embeddings have generally been classified in three broad categories: linear models, tensor decomposition models and neural models Hogan et al. (2021). Linear models use a distance to score triples, after applying a linear transformation to entity or relation embeddings, e.g. a translation Bordes et al. (2013) or rotation Sun et al. (2019). Region based models can essentially be viewed under the same umbrella, where the idea of computing distances between points is generalised to evaluating relationships between points and regions. Tensor decomposition models, such as ComplEx Trouillon et al. (2017); Lacroix et al. (2018), TuckER Balazevic et al. (2019b) and QuatE Zhang et al. (2019), view knowledge graphs as a three-dimensional tensor, which can be factorised into (low-rank) entity and relation embeddings. While these models tend to perform well, they often come with a significantly higher computational cost. Another important limitation is that they are far less interpretable in terms of how they capture semantic dependencies between relations. Bilinear models also have important theoretical limitations when it comes to capturing rules Gutiérrez-Basulto and Schockaert (2018). Neural models, including ConvE Dettmers et al. (2018) and graph neural network based approaches Zhu et al. (2021), are even less interpretable than tensor decomposition models.
@@
[…]
@@ 8 Conclusions
The study of region based knowledge graph embedding essentially aims to identify the simplest models that are capable of capturing particular types of reasoning patterns. Despite the extensive work on knowledge graph embedding, there are still surprisingly many open questions on this front. With the aim of furthering our understanding of region based models, we have studied an approach based on octagons. Octagons where shown to be capable of capturing arbitrary knowledge graphs, while a simpler model based on hexagons is not. Octagons also have the desired property of being closed under intersection and composition, which is not the case for closely related models such as BoxE and ExpressivE. We have furthermore shown that the octagon model is capable of modelling (particular) sets of composition rules, which is an important property that was not yet shown for existing region based models. Our empirical results show that octagon embeddings slightly underperform ExpressivE and further work is needed to better understand whether this performance gap can be closed by changing how the octagons are learned.
Acknowledgments
@@ B.1 Capturing Knowledge Graphs
Let us write $O^{s}_{r,e}$ for the octagon representing relation $r$ in coordinate associated with entity $e$ and relation $s$. For all coordinates where $s=r$, we choose $O^{s}_{r,e}=[0,2]\times[0,2]$. The octagons $O^{r}_{r,e}$ are chosen as follows:
@@
  • Now suppose $g=e$ with $e\neq f$. Then $e_{r,g}=e_{r,e}=0$. Since $(e,r,f)\in\mathcal{G}$, we have $f_{r,g}=1$. We thus have $(e_{r,g},f_{r,g})\in O^{r}_{r,g}$, regardless of whether $O^{r}_{r,g}=O^{+}$ or $O_{r,g}=O^{-}$.
@@
Now suppose $(e,r,f)\notin\mathcal{G}$. We need to show that there exists some $g\in\mathcal{E}$ such that $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$. If $e=f$ we know that $O^{r}_{r,e}=O^{-}$ since we assumed $(e,r,e)\notin\mathcal{G}$ Moreover, have $e_{r,e}=0$ and thus $(e_{r,e},e_{r,e})\notin O^{-}$. Now assume $e\neq f$. We show that $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$ for $g=e$. In other words, we show that $(e_{r,e},f_{r,e})\notin O^{r}_{r,e}$. We have $e_{r,e}=0$ and $f_{r,e}=2$. Regardless of whether $O^{r}_{r,g}=O^{+}$ or $O^{r}_{r,g}=O^{-}$ we thus find $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$. ∎
@@ B.2 Capturing Rules: Negative Results
Note that $\mathcal{K}$ is indeed consistent. In fact, we can capture the rules in $\mathcal{K}$ using a region-based embedding $\gamma$ in $\mathbb{R}^{3}$ as follows:
$\gamma(s_{1})$ =$=\{(0,0,0)\}$
$\gamma(s_{2})$ =$=\{(1,0,0)\}$
$\gamma(s_{3})$ =$=\{(0,1,0)\}$
$\gamma(s_{4})$ =$=\{(0,0,1)\}$
$\gamma(r_{1})$ $=\text{CH}(\{(1,0,0),(0,1,0),(0,0,1)\})$
$\gamma(r_{2})$ $=\text{CH}(\{(0,0,0),(0,1,0),(0,0,1)\})$
$\gamma(r_{3})$ $=\text{CH}(\{(0,0,0),(1,0,0),(0,0,1)\})$
$\gamma(r_{4})$ $=\text{CH}(\{(0,0,0),(1,0,0),(0,1,0)\})$

### v2

@@ 1 Introduction
To address these concerns, our aim in this paper is to develop a region based model which is as simple as possible, while (i) still being expressive enough to capture arbitrary knowledge graphs and (ii) using regions which are closed under intersection and composition. Based on these desiderata, we arrive at a model which relies on axis-aligned octagons (with all angles fixed at 45 degrees). We show that despite their simplicity, the proposed octagon embeddings are sufficiently expressive to properly capture a large class of rule bases. This is an important property, among others because it means that we can inject prior knowledge, in the form of a given rule base, without unintended consequences. Our result is considerably more general than what is possible with BoxE, which is not able to capture any closed path rules, and more general than what is known about ExpressivE. Moreover, because compositions and intersections of octagons can be straightforwardly computed, octagon embeddings are considerably more practical than existing alternatives. While our main focus is on better understanding the expressivity of region based models, we have also empirically evaluated the proposed octagon embeddings. We found octagons to achieve results close to the current state-of-the-art. This demonstrates that learning octagon embeddings is a promising strategy, especially in contexts where both knowledge graphs and rules need to be modelled. [^1] This paper is to appear in the proceedings of IJCAI 24.
@@ 2 Preliminaries
Figure 1: Basic region based embeddings: TransE bands (left) and hexagons (right).
@@ 4.2 Capturing Rules
Let us define the inverse region $X_{r}^{\textit{inv}}$ as follows:
$X_{r}^{\textit{inv}}{=}\{(x_{1},...,x_{n},y_{1},...,y_{n})\,|\,(y_{1},...,y_{n},x_{1},...,x_{n})\in X_{r}\}$
Note that when $X_{r}=[O_{1},\ldots,O_{n}]$ then $X_{r}^{\textit{inv}}=[O_{1}^{\textit{inv}},\allowbreak\ldots,\allowbreak O_{n}^{\textit{inv}}]$ with $O_{i}^{\textit{inv}}$ as defined in (9). Each of the aforementioned rule types can be straightforwardly modelled using regions. In particular, we say that a region based embedding $\gamma$ captures or satisfies (10) iff $\gamma(r)\subseteq\gamma(r)^{\textit{inv}}$; (11) iff $\gamma(r)\cap\gamma(r)^{\textit{inv}}=\emptyset$; (12) iff $\gamma(r_{1})\subseteq\gamma(r_{2})^{\textit{inv}}$; (13) iff $\gamma(r_{1})\subseteq\gamma(r_{2})$; (14) iff $\gamma(r_{1})\cap\gamma(r_{2})\subseteq\gamma(r_{3})$; (15) iff $\gamma(r_{1})\cap\gamma(r_{2})=\emptyset$; and (16) iff $\gamma(r_{1})\diamond\gamma(r_{2})\subseteq\gamma(r_{3})$. For a set of rules $\mathcal{K}$ and a rule $\rho$, we write $\mathcal{K}\models\rho$ to denote that $\rho$ is entailed by $\mathcal{K}$ in the usual sense. We also write $\gamma\models\rho$ to denote that the embedding $\gamma$ captures the rule $\rho$.
@@
This result is significant given the importance of composition rules for link prediction Meilicke et al. (2019), among others. Moreover, a similar result was not yet shown in previous work on region based embeddings, to the best of our knowledge. At the same time, however, the result is limited by the fact that only regular rules are considered and no inverse relations. Knowledge graph completion often relies (explicitly or implicitly) on rules such as $\textit{playsForTeam}(X,Y)\wedge\textit{playsForTeam}^{-1}(Y,Z)\wedge\textit{playsSport}(Z,U)\rightarrow\textit{playsSport}(X,U)$. Our analysis suggests that capturing sets of such rules would require an octagon model with cross-coordinate comparisons, which we leave as a topic for future work.
@@ 5 Learning Octagon Embeddings
where we write $\mathbf{u}_{c}=(\mathbf{u}^{+}+\mathbf{u}^{-})/2$ and $\mathbf{u}_{w}=(\mathbf{u}^{+}-\mathbf{u}^{-})/2$. The functions $\text{dist}_{x}$, $\text{dist}_{y}$ and $\text{dist}_{v}$ are defined analogously. Note that $\mathbf{e}\oplus\mathbf{f}\in X_{r}$ iff $\text{dist}_{x}(e,r,f)\geq 0.5$, $\text{dist}_{y}(e,r,f)\geq 0.5$, $\text{dist}_{u}(e,r,f)\geq 0.5$ and $\text{dist}_{v}(e,r,f)\geq 0.5$. Furthermore note that BoxE and ExpressivE also convert interval constraints into soft scores. However, rather than using scores of the form (17), these models use a piecewise linear function of the distance to the centre of the interval. Although a different slope is used for points inside and outside the interval, we may question to what extent the resulting embeddings still define regions. While the regions defined by scoring functions of the form (17) also have fuzzy boundaries, this “fuzziness” is mostly limited to the immediate vicinity of the boundary.
@@
where we write $\odot$ for component-wise multiplication. Weighted variants of $\text{dist}_{x}$, $\text{dist}_{y}$ and $\text{dist}_{v}$ are defined analogously. The attention weights intuitively allow the model to “forget” certain constraints, e.g. by focusing only (or primarily) on the $u$ band in some coordinates.
As in BoxE and ExpressivE, coordinate-wise scores are aggregated to a scalar by taking their negated norm. We define $\text{dist}(e,r,f)=\text{dist}_{x}(e,r,f)\oplus\text{dist}_{y}(e,r,f)\oplus\text{dist}_{u}(e,r,f)\oplus\text{dist}_{v}(e,r,f)$ and score triples as follows (and similar for the weighted variant):
$s(e,r,f)=-\|\>\text{dist}(e,r,f)\>\|_{p}$ (19)
In the above expression, $p$ is a hyperparameter. Learning occurs by minimizing the following margin loss function with self-adversarial negative sampling Sun et al. (2019):
@@
where $\mathbf{\alpha}=\text{softmax}(s(e_{1},r,f_{1})\oplus\ldots\oplus s(e_{n},r,f_{n})$, the margin $\lambda$ is a hyperparameter, and $N=\{(e_{1},r,f_{1}),\allowbreak\ldots,\allowbreak(e_{n},r,f_{n})\}$ is a set of randomly sampled negative examples.
@@ 6 Experimental Results
We consider several variants of our model. The full model is denoted by uvxy, whereas ux refers to a model in which $\text{dist}_{v}$ and $\text{dist}_{y}$ are not used, and similar for the other variants. For $uv$ and $uvxy$, we have used the variant with attention weights (18). For the other configurations, attention weights were not used, as they were found to make little difference [^4] An analysis of the impact of attention weights can be found in the appendix (Section C.2).. Note that our $u$ model is almost identical to TransE, except that an explicit width parameter is used, while uv can be seen as a variant of ExpressivE without learnable scale factors.
@@
$u$ 23.1 37.3 53.2 33.2 01.6 39.9 51.5 22.0
$ux$ 23.3 37.1 52.5 33.1 01.9 39.0 51.6 21.8
$uxy$ 23.2 37.2 53.1 33.2 01.8 40.8 52.8 22.8
$uv^{*}$ 24.1 36.9 52.8 33.6 43.6 48.5 52.9 46.9
$uvxy^{*}$ 24.1 36.7 51.7 33.2 43.6 49.2 56.1 47.9
Table 1: Link prediction performances of region based embedding models. Configurations with use the variant with attention weights.
@@
The performance of $u$ is almost identical to TransE, which suggests that the addition of an explicit width parameter plays no significant role. Adding the $x$, $y$ and $v$ constraints does not bring any benefits on FB15k-237. On WN18RR, however, introducing the $v$ constraint has a big impact on H@1, which goes from less than 1% for $u$, $ux$ and $uxy$ to more than 40% for $uv$ and $uvxy$. This result is in line with Proposition 5, which states that hexagons cannot capture relations that are both symmetric and irreflexive: for most WN18RR relations, models without $v$ constraints incorrectly rank $(e,r,e)$ triples first [^5] This result still holds without attention weights: on WN18RR, $uvxy$ achieves a H@1 of 39.6, H@3 of 45.9, H@10 of 50.7 and MRR of 43.8 without trainable attention weights.. Including the $x$ and $y$ constraints also improves performances in the case of WN18RR. Introducing attention weights, as in (18), has a significant positive effect on the overall performance on WN18RR but not for FB15k-237. As Table 1 shows, $uvxy$ outperforms BoxE on all metrics. Further analysis can be found in the appendix.
Octagon embeddings slightly underperform ExpressivE in most cases. We may wonder whether this reflects the inherent limitations of using fixed angles. In fact, when inspecting the scale factors in learned ExpressivE embedddings, we noticed that most are close to $-1$, 00 or $1$ (see Figure 4). In other words, most of the learned parallelograms are diamonds ($uv$), bands ($u$, $v$, $x$, $y$) or squares ($xy$), all of which can be represented in our octagon model. This suggests that the underperformance of the octagon model is rather the result of overparameterisation. On the one hand, we have shown theoretically that all four constraints are needed to capture semantic dependencies, and this is also confirmed by our empirical results on WN18RR. On the other hand, not all constraints might be needed for all coordinates. The use of attention weights was inspired by this view, but the weights themselves of course also introduce further parameters.
@@ 7 Related Work
Knowledge graph embeddings have generally been classified in three broad categories: linear models, tensor decomposition models and neural models Hogan et al. (2021). Linear models use a distance to score triples, after applying a linear transformation to entity or relation embeddings, e.g. a translation Bordes et al. (2013) or rotation Sun et al. (2019). Region based models can essentially be viewed under the same umbrella, where the idea of computing distances between points is generalised to evaluating relationships between points and regions. Tensor decomposition models, such as ComplEx Trouillon et al. (2017); Lacroix et al. (2018), TuckER Balazevic et al. (2019b) and QuatE Zhang et al. (2019), view knowledge graphs as a three-dimensional tensor, which can be factorised into (low-rank) entity and relation embeddings. While these models tend to perform well, they often come with a significantly higher computational cost. Another important limitation is that they are far less interpretable in terms of how they capture semantic dependencies between relations. Tensor decomposition models also have important theoretical limitations when it comes to capturing rules Gutiérrez-Basulto and Schockaert (2018). Neural models, including ConvE Dettmers et al. (2018) and graph neural network based approaches Zhu et al. (2021), are even less interpretable than tensor decomposition models.
@@
Region based approaches are closely related to ontology embeddings. In particular, several models have recently been introduced for the Description Logic $\mathcal{EL}^{++}$, which features subsumption and composition of relations, as well as union and intersection of concepts. Concepts are typically embedded as hyperballs Mondal et al. (2021) or hypercubes Xiong et al. (2022). The Box2EL model Jackermeier et al. (2023) also represents relations as regions, based on BoxE, which was found to improve performance. Given the limitations of BoxE on modeling composition, an evaluation of more expressive region based models is likely to yield even better results, although such an evaluation is yet to be performed.
@@ 8 Conclusions
The study of region based knowledge graph embedding essentially aims to identify the simplest models that are capable of capturing particular types of reasoning patterns. Despite the extensive work on knowledge graph embedding, there are still surprisingly many open questions on this front. Octagons where shown to be capable of capturing arbitrary knowledge graphs, while a simpler model based on hexagons is not. Octagons also have the desired property of being closed under intersection and composition, which is not the case for closely related models such as BoxE and ExpressivE. We have furthermore shown that the octagon model is capable of modelling (particular) sets of composition rules, which is an important property that was not yet shown for existing region based models. Our empirical results show that octagon embeddings slightly underperform ExpressivE and further work is needed to better understand whether this performance gap can be closed by changing how the octagons are learned.
@@ B.1 Capturing Knowledge Graphs
Let us write $O^{s}_{r,e}$ for the octagon representing relation $r$ in coordinate associated with entity $e$ and relation $s$. For all coordinates where $s\neq r$, we choose $O^{s}_{r,e}=[0,2]\times[0,2]$. The octagons $O^{r}_{r,e}$ are chosen as follows:
@@
  • Now suppose $g=e$ with $e\neq f$. Then $e_{r,g}=e_{r,e}=0$. Since $(e,r,f)\in\mathcal{G}$, we have $f_{r,g}=1$. We thus have $(e_{r,g},f_{r,g})\in O^{r}_{r,g}$, regardless of whether $O^{r}_{r,g}=O^{+}$ or $O^{r}_{r,g}=O^{-}$.
@@
Now suppose $(e,r,f)\notin\mathcal{G}$. We need to show that there exists some $g\in\mathcal{E}$ such that $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$. If $e=f$ we know that $O^{r}_{r,e}=O^{-}$ since we assumed $(e,r,e)\notin\mathcal{G}$ Moreover, we have $e_{r,e}=0$ and thus $(e_{r,e},e_{r,e})\notin O^{-}$. Now assume $e\neq f$. We show that $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$ for $g=e$. In other words, we show that $(e_{r,e},f_{r,e})\notin O^{r}_{r,e}$. We have $e_{r,e}=0$ and $f_{r,e}=2$. Regardless of whether $O^{r}_{r,g}=O^{+}$ or $O^{r}_{r,g}=O^{-}$ we thus find $(e_{r,g},f_{r,g})\notin O^{r}_{r,g}$. ∎
@@ B.2 Capturing Rules: Negative Results
Note that $\mathcal{K}$ is indeed consistent. In fact, we can capture the rules in $\mathcal{K}$ using a region-based embedding $\gamma$ in $\mathbb{R}^{4}$ as follows:
$\gamma(s_{1})$ =$=\{(0,0,0,0)\}$
$\gamma(s_{2})$ =$=\{(1,0,0,0)\}$
$\gamma(s_{3})$ =$=\{(0,1,0,0)\}$
$\gamma(s_{4})$ =$=\{(0,0,1,0)\}$
$\gamma(r_{1})$ $=\text{CH}(\{(1,0,0,0),(0,1,0,0),(0,0,1,0)\})$
$\gamma(r_{2})$ $=\text{CH}(\{(0,0,0,0),(0,1,0,0),(0,0,1,0)\})$
$\gamma(r_{3})$ $=\text{CH}(\{(0,0,0,0),(1,0,0,0),(0,0,1,0)\})$
$\gamma(r_{4})$ $=\text{CH}(\{(0,0,0,0),(1,0,0,0),(0,1,0,0)\})$